OpenStudy (anonymous):

Which of the following is an even function? g(x) = (x – 1)2 + 1 g(x) = 2x2 + 1 g(x) = 4x + 2 g(x) = 2x

OpenStudy (anonymous):

@Lokid

OpenStudy (anonymous):

@Kash_TheSmartGuy

OpenStudy (anonymous):

sorry i don't rember :/

OpenStudy (anonymous):

f(-a)=f(a) then f(x) is an even function

OpenStudy (anonymous):

listen to that person ^

OpenStudy (anonymous):

@fakeee I'm confused

OpenStudy (anonymous):

I am saying that f(x) is an even function, if it satisfies f(a)=f(-a) for any value of a. for example, you have to have f(1)=f(-1)

OpenStudy (anonymous):

in your options, there is only one such function that satisfies f(-a)=f(a) which on do you think it is and why?

OpenStudy (anonymous):

So it's A. g(x) = (x – 1)2 + 1 ?

OpenStudy (anonymous):

plug in a and -a into each option and set them equivalent to each other: (like this ) \((a-1)^2+1=(-a-1)^2+1\) \(2(a)^2+1=2(-a)^2+1\) \(4(a)+2=4(-a)+2\) \(2(a)=2(-a)\)

OpenStudy (anonymous):

can you tell me now which is that function that satisfies f(a)=f(-a) ?

OpenStudy (anonymous):

So I plug in a for x ?

OpenStudy (anonymous):

Nvm I figured it out thanks you